OpenStudy (mpj4):

The two sides of a triangle are 17 cm and 28 cm long, and the length of the median drawn to the third side is equal to 19.5 cm. Find the distance from an endpoint of this median to the longest side What does "find the distance from an endpoint of this median to the longest side" mean? I interpreted it as making a perpendicular line from the longest side to the endpoint of said median.

OpenStudy (mpj4):

 There's no drawing so this is just a drawing of what i think it looks like.

OpenStudy (mpj4):

nvm, I found the answer.. I think. 19.5=1/2√2a^2+2b^2-c^2 39=√2(28)^2+2(17)^2-c^2 39=√1568+578-c^2 39=√2146-c^2 39^2=2146-c^2 1521=2146-c^2 c^2=2146-1521 c^2=625 √c^2=√625 c=25  This is for all those solid mensuration students that couldn't figure this one out, if they can still find this xD